media%2Ff34%2Ff3499ba2-430b-4f2a-ac59-bc
Answer #1

The uncertainity can be given as,

\Delta x*\Delta p\geq \frac{h}{4\pi }

Where delta x is the uncertainity in position, delta p is the uncertainity in momemtum and h is the Planck constant.

Momentum can be written as,

\Delta x*m\Delta v\geq \frac{h}{4\pi }

where, m is the mass and delta v is the uncertainity in velocity.

Now, we have to substitute the values.

\Delta x*0.230kg*0.1*0.277778m/s\geq \frac{h}{4\pi }

(be cautious of the units)

\Delta x\geq \frac{h}{4\pi *0.230kg*0.1*0.277778m/s}

\Delta x\geq 8.253*10^{-33}m